OpenStudy (anonymous):

Find the inverse of the function F(x) = 5+ln(x+1)

OpenStudy (anonymous):

The inverse of a function is the function that undoes its operation. The notation for the inverse of the function f is f-1. In function notation this can be written like this. f-1(f(x))=x f(f-1(x))=x

OpenStudy (anonymous):

y = 5/(x + 1) Finding the inverse of ANY function f(x) requires these steps: 1) Change f(x) into y (if it's not changed already). 2) Interchange the x and y variables. 3) Solve for y. The result of y will be your function. 4) Change the y to f inverse of x, or f^(-1)(x). Let's assume you started off with f(x) = 5/(x + 1) Change f(x) into y. y = 5/(x + 1) Interchange x and y. x = 5/(y + 1) Solve for y. x(y + 1) = 5 y + 1 = 5/x y = (5/x) - 1 Change y into f^(-1)(x) (concluding statement) f^(-1)(x) = (5/x) - 1 This is the inverse of f(x) = 5/(x + 1)